Fermi National Laboratory
 

Operation Summary for April 30, 2010 - May 3, 2010

 

The Integrated Luminosity for the period from 4/26/10 to 5/3/10 was 70.55 pb-1.
NuMI reported receiving 8.46E18 protons on target during this same period.

Friday April 30

The day shift began with Operations monitoring the Tevatron (TeV) store 7778 (118.27E30) and the Recycler stash (296.7E10), with the Antiproton Source (Pbar) stacking (24.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI, MiniBooNE, and Meson MTest experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, Operations terminated store 7778 at 1 PM, put the TeV into HEP shot setup, and established store 7779 at 3:10 PM with an average initial luminosity (AIL) of 373.0E30. Operations prepared the Pbar rings for access at 1 PM to search for a LCW leak. Water group personnel fixed the leak and Pbar resumed stacking at 3 PM.

During the evening shift, Operations disabled beam to MTest at 6 PM and MI personnel conducted slow-spill studies until 7:45 PM.

Saturday May 1

The midnight shift began with Operations monitoring store 7779 (116.41E30) and the stash (208.2E10), with Pbar stacking (27.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to MTest at 4:20 AM. A Linac RF station (KRF3) tripped off three times; it reset.

During the day shift, Operations terminated store 7779 at 9 AM, put the TeV into shot setup, and established store 7780 at 10:15 AM with an AIL of 383.0E30.

During the evening shift, Operations disabled the MTest beam around 7 PM.

Sunday May 2

The midnight shift began with Operations monitoring store 7780 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Linac quadrupole power supply tripped off and required a control card replacement. Operations terminated store 7780 at 3:10 AM and put the TeV into shot setup. Unfortunately, the TeV quenched at low beta sequence 23. The investigation did not reveal a cause. Operators noticed that proton positions during the squeeze were not normal. Operations reported that a TeV sector C48 Kautzky valve would not close, so the switch off of TeV did not begin until the end of the shift. Also, at the end of shift, Mechanical Support vacuum technicians were en route to repair the Tevatron Electron Lens (TEL2) vacuum. Operations established beam to MTest after the quench and disabled it at the end of shift due to the pending Transfer Hall access.

During the day shift, preparations for the TeV A-E access and Transfer Hall access continued. The Kautzky valve at C48 closed prior to accessing the TeV; CHL personnel accessed anyway to make sure there wasn't an ice ball. Mechanical support personnel accessed to repair the TEL2 vacuum and were out of the tunnel by 11:30 AM. Operations prepared the TeV for aperture scans at 2:30 PM, which continued until 6 PM. Operators conducted orbit checks and then put the TeV into shot setup at 7 PM. The NuMI extraction kicker tripped off during shot setup and held off beam.

During the evening shift, Operations established store 7784 at 8:30 PM with an AIL of 355.0E30. NuMI kicker personnel investigated and determined an arc occurred internal to the kicker power supply. They pulled electronics from the oil tank and found the source of the arc was in a connector assembly.

Monday May 3

The midnight shift began with Operations monitoring store 7784 and the stash, with Pbar stacking, with NuMI off due to a power supply problem, and with MiniBooNE taking beam.

During the midnight shift, NuMI personnel completed their kicker repairs. Operations resumed sending beam to NuMI at 4 AM. The kicker started causing problems again; Operations disabled the NuMI beam at 6:20 AM.

The Plans for Monday

The plans for today are to stack, stash, and store, and to repair the NuMI kicker power supply. The Run Coordinator will allow 90 minutes for the following work this afternoon after store 7784 is terminated: D0 and CDF will work on their detector, cryo system personnel will conduct wet engine maintenance, and MI personnel will check their coalescing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/03/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory